# Scaling of Disorder Operator and Entanglement Entropy at Easy-Plane Deconfined Quantum Criticalities

### v2

@@ Abstract
We systematically investigate the scaling behavior of the disorder operator and the entanglement entropy (EE) of the easy-plane JQ (EPJQ) model at its transitions between the antiferromagnetic XY ordered phase (AFXY) and the valence bond solid (VBS) phase. We find (1) there exists a tiny yet finite value of the order parameters at the AFXY-VBS phase transition points of the EPJQ model, and the finite order parameter is strengthened as anisotropy $\Delta$ varies from the Heisenberg limit ($\Delta=1$) to the easy-plane limit ($\Delta=0$); (2) Both EE and disorder operator with smooth boundary cut exhibit anomalous scaling behavior at the transition points, resembling the scaling inside the Goldstone model phase, and the anomalous scaling becomes strengthened as the transition becomes more first order; (3) First put forward in Ref. [arXiv:2401.12838], with the finite-size corrections in EE for Goldstone phase is properly considered in the fitting form, the anomalous scaling behavior of EE can be adapted with emergent SO(5) symmetry breaking at the Heisenberg limit ($\Delta=1$). We extend this method in the EPJQ model and observe similar yet weaker results, which may indicate emergent SO(4) symmetry breaking in the easy-plane regime ($\Delta<1$) or emergent SO(5) symmetry breaking in the Heisenberg limit ($\Delta=1$). These observations provide evidence that the Néel-VBS transition in the JQ model setting evolves from weak to prominent first-order transition as the system becomes anisotropic, and the non-local probes such as EE and disorder operator, serve as the sensitive tool to detect such salient yet fundamental features.
@@ I Introduction
The conventional Landau-Ginzburg-Wilson (LGW) paradigm for characterizing phases and their transitions has been confronted with the notion of deconfined quantum critical points (DQCPs), which describes a direct continuous transition between Néel ordered phases and valence-bond-solid (VBS) phases Senthil (2004); Senthil et al. (2004); Levin and Senthil (2004); Senthil et al. (2005). However, the feasibility of realizing such transitions in lattice models remains a contentious issue because of the observations of various anomalous behavior against conformal field theories (CFT) Harada et al. (2013); Chen et al. (2013); Nahum et al. (2015); Nakayama and Ohtsuki (2016); Li (2022); Poland et al. (2019); Wang et al. (2022); Zhao et al. (2022a); Song et al. (2023a, b); Liu et al. (2023); Liao et al. (2023); Liu et al. (2024), such as drifting of critical exponents Harada et al. (2013); Chen et al. (2013); Nahum et al. (2015), violation of CFT bounds Nakayama and Ohtsuki (2016); Li (2022); Poland et al. (2019), and the anomalous scaling behavior of entanglement entropies and disorder operators Wang et al. (2022); Zhao et al. (2022a); Song et al. (2023a, b); Liu et al. (2023); Liao et al. (2023); Liu et al. (2024). A deeper exploration of the origins of these anomalous behaviors will benefit not only the better understanding of nature of DQCPs but also its lattice model and even material realizations Zayed et al. (2017); Guo et al. (2020); Jiménez et al. (2021); Sun et al. (2021); Cui et al. (2023); Guo et al. (2023).
@@
Among all the anomalous behavior against CFTs, we are in particular interested in the anomalous finite-size scaling behavior of Rényi entanglement entropy (EE) and disorder operator, as it happens even at relatively small system sizes Wang et al. (2022); Zhao et al. (2022a); Song et al. (2023a, b); Liu et al. (2023); Liao et al. (2023); Liu et al. (2024) where other physical quantities such as order parameters and correlation functions still exhibit good agreement with a continuous phase transition Sandvik (2007); Lou et al. (2009); Shao et al. (2016); Ma et al. (2018a); Qin et al. (2017). In fact, at the DQCPs of SU(2) JQ2 and JQ3 models, our previous works Zhao et al. (2022a); Song et al. (2023a); Wang et al. (2022); Song et al. (2023b) show that disorder operator and EE all exhibit anomalous scaling against CFTs. To be specific, for the corner cut case where the boundary of subregion has four $\frac{\pi}{2}$ corners, both disorder operator and EE have a positive logarithmic correction to the leading order area law term which violates the unitary CFT prediction that the correction must be negative  Casini and Huerta (2012); Fradkin and Moore (2006); Casini and Huerta (2007). Later, further studies clarified that the violation of EE actually arises from the smooth part of the subregion. In fact, for smooth cut case where the boundary has no sharp corners, the anomalous logarithmic subleading correction also exists in EE, which is incompatible with CFTs Song et al. (2023b, a), but seems to be more compatible with the existence of Goldstone modes  Metlitski and Grover (2015); Deng et al. (2024) and can also possibly be adapted with the ‘walking" pseudo-criticality behavior at the transition Wang et al. (2017); Nahum (2020); Ma and Wang (2020); Zhou et al. (2023). In addition, recently another work D’Emidio and Sandvik (2024) shows that scaling of EE at the DQCP is cut-dependent, and the anomalous log correction for smooth cut can be suppressed by considering a tilted square lattice by $\frac{\pi}{4}$ degrees, and the essential information of the CFTs might be captured by this cut at small system sizes. In this context, the scalings of these nonlocal physical observables prove to be powerful and sensitive tools for diagnosing the various possible scenarios of the DQCPs Zhao et al. (2022a); Song et al. (2023a); Wang et al. (2022); Song et al. (2023b); Deng et al. (2024) and revealing the universal information of the possible CFT nearby the transitions D’Emidio and Sandvik (2024).
@@
Here in this work, we aim to shed more insights on this issue by tunning the JQ-type Néel-VBS transition from weakly first-order to a more prominent first-order transition, such that many of the salient features observed at the SU(2) JQ limit, for example, the anomalous finite-size scaling behavior in the entanglement entropy (EE) Zhao et al. (2022a); Song et al. (2023a, b); Deng et al. (2024) and disorder operator Wang et al. (2022), can also be examined at a clear first-order transition, and the connection between the observed anomalous behavior in EE and disorder operator with the first-order nature of the transitions can be further clarified. A possible platform to achieve this goal is the easy-plane JQ (EPJQ) model, which adds the easy-plane anisotropy to the Heisenberg term in the JQ model, as shown in Eq. (1), to tune the Néel-VBS transition to be more first-ordered. Previous studies have investigated the nature of the Néel-VBS transition at different anisotropy $\Delta$, both in 2D and 3D lattice model settings Qin et al. (2017); Ma et al. (2019); Zhao et al. (2019); Desai and Kaul (2020); Sun et al. (2021), and the general expectation/observation is that the more $\Delta$ tunes away from the Heisenberg limit ($\Delta=1$) towards the easy-plane limit ($\Delta=0$), the stronger the first order of the Néel-VBS transition. The EPJQ model thus serves as an ideal platform to study the anomalous scaling behavior of EE and disorder operator at these transition points simply by tuning the values of $\Delta$.
@@
In this paper, we use the stochastic series expansion (SSE) quantum Monte Carlo (QMC) methods Sandvik (1999); Syljuåsen and Sandvik (2002) to systematically study the order parameter, entanglement entropy, and disorder operator at the transitions of the EPJQ model. We carefully analyze the remaining ordered moments at the transitions and systematically study the evolution of the scaling of entanglement entropy and disorder operator along the critical line as a function of $\Delta$ in Eq. (1). Our major findings are
@@
    For $\Delta\in[0,1)$ there always exist tiny yet finite order parameters at the AFMXY/Néel-VBS phase transitions of the model (including the finite order parameter at the $\Delta=1$ limit Takahashi et al. (2024)), indicating the first-order nature of the transition, and the moment increases as $\Delta$ is tuned from 1 to 0;
@@
    Both EE and disorder operator with standard smooth boundary cut (without tilting) exhibit anomalous scaling behavior against CFTs at the transition points, and resemble more to the scalings inside the Goldstone model phase, and the anomalous scalings also become strengthened as the transition becomes more first order;
@@
    Pioneered in Ref. Deng et al. (2023a, 2024), with the finite-size corrections in EE for Goldstone phase is properly dealt with, the anomalous scaling behavior of EE can be adapted with emergent SO(5) symmetry breaking at the Heisenberg limit ($\Delta=1$). We extend this method in EPJQ model and observe similar yet weaker results which may indicate emergent SO(4) symmetry breaking in the easy-plane regime ($\Delta<1$) or emergent SO(5) symmetry breaking in the Heisenberg limit ($\Delta=1$).
@@
Our work thus provides a comprehensive study of the scaling of EE and disorder operator along the critical line where the weakly first-order transition is gradually tuned to a prominent first-order one by tunning $\Delta=1$ to $\Delta=0$. We show that the anomalous scalings of EE and disorder operator are positively related with the first order behavior, proving a strong evidence that the DQCP of JQ model is indeed a very weakly-first-order transition and the anomalous log-corrections come from Goldstone modes. Apart from this, our work supports the findings in Ref. Deng et al. (2024) that the origin of anomalous scaling of EE at the transition of $\Delta=1$ is the emergent continuous SO(5) symmetry breaking and the Goldstone modes associated with it. However, our results are weaker as we do not observe the same good convergence behavior to exact number of Goldstone modes required by the SO(5) symmetry breaking at $\Delta=1$ or SO(4) symmetry breaking at $\Delta<1$, which might originate from insufficient data quality or the less pronounced emergent SO(4) symmetry behavior in the easy-plane case. Despite that, our results strongly suggest that the DQCPs of EPJQ model are first-ordered for all $\Delta$ and scalings of EE and disorder operator serve as sensitive tools for detecting such transitions.
@@
The rest of the paper is organized as follows. In Sec. II, we first introduce the EPJQ model and discuss how we extract the finite order parameters at its transition points, then we move on to the discussion of disorder operator in Sec. III and reveal its anomalous scaling with comparisons to that of the conventional (2+1)D O(3) QCP and Néel phase, next we discuss the results of EE at the EPJQ transitions in Sec. IV and point out its anomalous scaling very likely stems from the residual Goldstone mode at the first-order transition points, finally we give a comprehensive summary on our results, discuss the relation of our work with other recent works Deng et al. (2024); D’Emidio and Sandvik (2024) and point out future directions in Sec. V. Detailed analysis of the determination of the transition points and extrapolations of the order parameter, and the analysis of the quality of the fitting in EE data are presented in the Supplemental Material (SM) sup .
@@
Figure 1: (a) Ground state phase diagram of the easy-plane JQ3 model. The red dots are critical points $q_{c}$ determined from FSS analysis of Binder cumulants for different $\Delta$ values. When $\Delta=1$, the model is the standard JQ3 model and the weakly-first-order transition point Takahashi et al. (2024) separates the Néel order state with the VBS state. At finite $\Delta$, the antiferromagnetic Néel order is of the XY type. The Néel-VBS transition becomes prominently first order as $\Delta$ goes from 1 to 0. The insets exhibit the EPJQ model (left) and the entanglement region with smooth boundaries (right). (b) The non-zero order parameters for both AFXY and VBS phases at $q_{c}(\infty)$ with $\Delta$ changes from 00 to $0.6$. At $\Delta=1$, the extrapolated order parameters are from Ref. Takahashi et al. (2024) of JQ2 model.
II Model and Phase Diagram
@@
We study the easy-plane JQ3 (EPJQ) model Qin et al. (2017); Ma et al. (2019); Zhao et al. (2019); Desai and Kaul (2020); Sun et al. (2021) as illustrated in the lower left inset of Fig. 1 (a) with the following Hamiltonian
@@
where $P_{ij}=\frac{1}{4}-\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ is the two-spin singlet projector. For $\Delta=1$, the model reduces to JQ3 model with isotropic Heisenberg interactions. It has been found a direct quantum phase transition in the JQ3 model between the Néel and VBS phases happens at $q_{c}=0.59864(5)$ (with $q=\frac{Q}{J+Q}$ and $J+Q=1$ as the energy unit) Lou et al. (2009); Wang et al. (2022). However, more numerical evidence, especially from the scaling of nonlocal observables such as EE Zhao et al. (2022a); Song et al. (2023a, b) and disorder operator Wang et al. (2022), reveal an anomalous behavior against CFTs even if the partitioning of the lattice is smooth, i.e. no sharp corners on the boundary. The fact that the sign of the observed log-coefficient is consistent with that of the presence of the Goldstone mode, further suggests there exist finite antiferromagnetic moment, i.e. remaining of the Néel order, at the transition point Metlitski and Grover (2015). Such evidence promotes the understanding that the Néel-VBS transition is indeed first order, despite being very weak at the SU(2) limit Takahashi et al. (2024). In this work, we monitor the behavior of both conventional order parameters and nonlocal observables at such transitions, as the anisotropy $\Delta$ in Eq. (1) is tuned from the Heisenberg limit($\Delta=1$) to easy-plane limit ($\Delta=0$).
@@
Table 1: The critical points obtained from the Binder cumulants of $U_{M}$ and $U_{D}$ as shown in Fig. 2 for different $\Delta=0,0.1,0.3,0.4,0.5$ and 0.6 using similar analysis in Fig. 2. The $q_{c}(\infty)$ of $U_{M}$ and $U_{D}$ are identical within two sigma. The detailed analyses are shown in Sec. I of SM sup .
At each $\Delta\in[0,1)$, the phase transition between the antiferromagnetic XY (AFXY) and VBS phase at zero temperature happens when tunning $q=Q/(J+Q)$ from zero to $q_{c}$. To determine the critical points, we perform SSE-QMC simulations Sandvik (1999); Syljuåsen and Sandvik (2002) on the EPJQ model at $\beta=2L$ for different system sizes $L=16,24,32,48,64,96,128$ with $\Delta=0$, $0.1$, $0.3$, $0.5$ and $0.6$. In the AFXY phase, the order parameter is defined as the sublattice magnetization which breaks the U(1) symmetry, with the $x$ component being written as
@@
where $\{x_{i}$,$y_{i}\}$ are coordinates of site $i$ in the lattice. In the simulation with finite system sizes, the expectation value of $\langle M_{x}^{2}\rangle$ can be viewed as the square of the AFXY order parameter, whereas $\langle M_{x}\rangle=0$ at finite sizes.
@@
We thus perform the finite-size scaling (FSS) analysis of $\langle M^{2}_{x}\rangle$ and $\langle D^{2}\rangle$ to determine the transition points at the thermodynamic limit (TDL) Ma et al. (2018b). The dimensionless quantity Binder cumulants are widely used in the FSS to extract the critical points. For those two ground states in our study, the Binder cumulants are defined as
@@
with $\nu$ the correlation length exponent and $\omega$ the finite-size correction exponent Qin et al. (2017); Ma et al. (2018b); Chen et al. (2023).
@@
Repeating such a procedure, we can also get $q_{c}(\infty)$ for different $\Delta$ in Tab. 1. For a given $\Delta<1$, the $q_{c}(\infty)$ obtained from crossings of different dimensionless quantities agree with each other considering the error as large as two sigma, which can be true for the continuous phase transition as well as the first-order one. However, we should point out that $U_{M}$ of the AFXY order parameter for $\Delta=0.1$ becomes negative close to $q_{c}$ at the largest simulated size $L=128$, as shown in Fig. 2 (a), the negative $U_{M}$ can also be regarded as a signature of the first-order phase transition Binder and Landau (1984).
We have performed the extrapolation of the ordered moments as in Fig. 2 (d) for all the $\Delta$ values, and the obtained results are summarized in Fig. 1 (b). It is clear that $\langle M^{2}_{x}\rangle$ and $\langle D^{2}\rangle$ are small at $\Delta=0.6$ (they are even smaller at the isotropic limit), but they gradually increases as $\Delta\to 0$. We show all these extrapolations in the Sec. I of SM sup . The enhancement of the ordered moments as $\Delta\to 0$ along the phase AFXY-VBS phase boundary in Fig. 1 (a) suggests evolution from very weakly first-order to stronger ones. As will be shown below, our non-local EE and disorder operator measurements (with smaller systems sizes compared with order parameters) can also capture such intriguing features.
@@
where $\langle ij\rangle$ denotes the $J_{1}$ bond and $\langle ij\rangle^{{}^{\prime}}$ denotes the $J_{2}$ bond. This model has a well-established (2+1)d O(3) QCP at $(J_{2}/J_{1})_{c}=1.90951(1)$ Ma et al. (2018b) separating the Néel phase and a symmetric singlet product phase. As will also shown below, the scaling of disorder operator of this O(3) QCP does not have anomalous log-corrections.

### v3

@@ Abstract
We systematically investigate the scaling behaviors of the disorder operator and the entanglement entropy (EE) of the easy-plane JQ (EPJQ) model at its transitions between the antiferromagnetic XY ordered phase (AFXY) and the valence bond solid (VBS) phase. We find there exists a tiny yet finite value of the order parameters at the AFXY-VBS phase transition points of the EPJQ model, and the finite order parameter is strengthened as anisotropy $\Delta$ varies from the Heisenberg limit ($\Delta=1$) to the easy-plane limit ($\Delta=0$). This observation provides evidence that the Néel-VBS transition in the JQ model setting evolves from weak to prominent first-order transition as the system becomes anisotropic. Furthermore, both EE and disorder operator with smooth boundary cut exhibit anomalous scaling behavior at the transition points, resembling the scaling inside the Goldstone mode (AFXY) phase, and the anomalous scaling becomes strengthened as the transition becomes more first order. In particular, for $\Delta\leq 0.3$, the obtained log-coefficients converge to 0.5 which is the same as the contribution from one Goldstone mode in the Néel phase. For $\Delta>0.3$, the log-coefficients are smaller and our findings might suffer from strong finite-size effects due to the fact that the remaining Néel order here is quite tiny.
@@ I Introduction
The conventional Landau-Ginzburg-Wilson (LGW) paradigm for characterizing phases and their transitions has been confronted with the notion of deconfined quantum critical points (DQCPs), which describes a direct continuous transition between Néel ordered phases and valence-bond-solid (VBS) phases [1, 2, 3, 4]. However, the feasibility of realizing such transitions in lattice models remains a contentious issue because of the observations of various anomalous behavior against conformal field theories (CFT) [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], such as drifting of critical exponents [5, 6, 7], violation of CFT bounds [8, 9, 10], and the anomalous scaling behavior of entanglement entropies and disorder operators [11, 12, 13, 14, 15, 16, 17]. A deeper exploration of the origins of these anomalous behaviors will benefit not only the better understanding of nature of DQCPs but also its lattice model and even material realizations [18, 19, 20, 21, 22, 23].
@@
Among all the anomalous behavior against CFTs, we are in particular interested in the anomalous finite-size scaling behavior of Rényi entanglement entropy (EE) and disorder operator, as it happens even at relatively small system sizes [11, 12, 13, 14, 15, 16, 17] where other physical quantities such as order parameters and correlation functions still exhibit good agreement with a continuous phase transition [24, 25, 26, 27, 28]. In fact, at the DQCPs of SU(2) JQ2 and JQ3 models, our previous works [12, 13, 11, 14] show that disorder operator and EE all exhibit anomalous scaling against CFTs. To be specific, for the corner cut case where the boundary of subregion has four $\frac{\pi}{2}$ corners, both disorder operator and EE have a positive logarithmic correction to the leading order area law term which violates the unitary CFT prediction that the correction must be negative  [29, 30, 31]. Later, further studies clarified that the violation of EE actually arises from the smooth part of the subregion. In fact, for smooth cut case where the boundary has no sharp corners, the anomalous logarithmic subleading correction also exists in EE, which is incompatible with CFTs [14, 13], but seems to be more compatible with the existence of Goldstone modes  [32, 33] and can also possibly be adapted with the ‘walking" pseudo-criticality behavior at the transition [34, 35, 36, 37]. In addition, recently another work [38] shows that scaling of EE at the DQCP is cut-dependent, and the anomalous log correction for smooth cut can be suppressed by considering a tilted square lattice by $\frac{\pi}{4}$ degrees, and the essential information of the CFTs might be captured by this cut at small system sizes. In this context, the scalings of these nonlocal physical observables prove to be powerful and sensitive tools for diagnosing the various possible scenarios of the DQCPs [12, 13, 11, 14, 33] and revealing the universal information of the possible CFT nearby the transitions [38].
@@
Here in this work, we aim to shed more insights on this issue by tunning the JQ-type Néel-VBS transition from weakly first-order [39] to a more prominent first-order transition, such that many of the salient features observed at the SU(2) JQ limit, for example, the anomalous finite-size scaling behavior in the entanglement entropy (EE) [12, 13, 14, 33] and disorder operator [11], can also be examined at a clear first-order transition, and the connection between the observed anomalous behavior in EE and disorder operator with the first-order nature of the transitions can be further clarified. A possible platform to achieve this goal is the easy-plane JQ (EPJQ) model, which adds the easy-plane anisotropy to the Heisenberg term in the JQ model, as shown in Eq. (1), to tune the Néel-VBS transition to be more first-ordered. Previous studies have investigated the nature of the Néel-VBS transition at different anisotropy $\Delta$, both in 2D and 3D lattice model settings [28, 40, 41, 42, 21], and the general expectation/observation is that the more $\Delta$ tunes away from the Heisenberg limit ($\Delta=1$) towards the easy-plane limit ($\Delta=0$), the stronger the first order of the Néel-VBS transition. The EPJQ model thus serves as an ideal platform to study the anomalous scaling behavior of EE and disorder operator at these transition points simply by tuning the values of $\Delta$.
@@
In this paper, we use the stochastic series expansion (SSE) quantum Monte Carlo (QMC) methods [43, 44] to systematically study the order parameter, entanglement entropy, and disorder operator at the transitions of the EPJQ model. We carefully analyze the remaining ordered moments at the transitions and systematically study the evolution of the scaling of entanglement entropy and disorder operator along the critical line as a function of $\Delta$ in Eq. (1). Our major findings are
@@
    For $\Delta\in[0,1)$ there always exist tiny yet finite order parameters at the AFMXY/Néel-VBS phase transitions of the model (including the finite order parameter at the $\Delta=1$ limit [39]), indicating the first-order nature of the transition, and the moment increases as $\Delta$ is tuned from 1 to 0;
@@
    Both EE and disorder operator with standard smooth boundary cut (without tilting) exhibit anomalous scaling behavior against CFTs at the transition points, which resemble the the scalings inside the Goldstone model phase, and the anomalous scalings also become strengthened as the transition becomes more first order.
@@
    For $\Delta\leq 0.3$, the obtained log-coefficients converge to 0.5 which is the same as the contribution from one Goldstone mode in the Néel phase. For $\Delta>0.3$, the log-coefficients are smaller than 0.5, and that might be caused by strong finite-size effects.
@@
Our work thus provides a comprehensive study of the scaling of EE and disorder operator along the critical line where the weakly first-order transition is gradually tuned to a prominent first-order one by tunning $\Delta=1$ to $\Delta=0$. We show that the anomalous scalings of EE and disorder operator are very positively related with the first order behavior,suggesting that the DQCP of JQ model is indeed a very weakly-first-order transition and the anomalous log-corrections come from Goldstone modes.
@@
The rest of the paper is organized as follows. We first introduce the EPJQ model and discuss how we extract the finite order parameters at its transition points, then we move on to the discussion of disorder operator and reveal its anomalous scaling with comparisons to that of the conventional (2+1)D O(3) QCP and Néel phase, next we discuss the results of EE at the EPJQ transitions and point out its anomalous scaling very likely stems from the residual Goldstone mode at the first-order transition points, finally we give a comprehensive summary on our results and point out future directions. Detailed analysis of the determination of the transition points and extrapolations of the order parameter, and the analysis of the quality of the fitting in EE data are presented in the Appendix  B.
@@
Figure 1: Phase diagram of the EPJQ model and finite order parameter at the transitions. (a) Ground state phase diagram of the easy-plane JQ3 model. The red dots are critical points $q_{c}$ determined from FSS analysis of Binder cumulants for different $\Delta$ values. When $\Delta=1$, the model is the standard JQ3 model and the weakly-first-order transition point [39] separates the Néel order state with the VBS state. At finite $\Delta$, the antiferromagnetic Néel order is of the XY type. The Néel-VBS transition becomes prominently first order as $\Delta$ goes from 1 to 0. The insets exhibit the EPJQ model (left) and the entanglement region with smooth boundaries (right). (b) The non-zero order parameters for both AFXY and VBS phases at $q_{c}(\infty)$ with $\Delta$ changes from 0 to $0.6$. At $\Delta=1$, the extrapolated order parameters are from Ref. [39] of JQ2 model.
II Model and Phase Diagram.
@@
We study the easy-plane JQ3 (EPJQ) model [28, 40, 41, 42, 21] as illustrated in the lower left inset of Fig. 1 (a) with the following Hamiltonian
@@
where $P_{ij}=\frac{1}{4}-\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ is the two-spin singlet projector. For $\Delta=1$, the model reduces to JQ3 model with isotropic Heisenberg interactions. It has been found a direct quantum phase transition in the JQ3 model between the Néel and VBS phases happens at $q_{c}=0.59864(5)$ (with $q=\frac{Q}{J+Q}$ and $J+Q=1$ as the energy unit) [25, 11]. However, more numerical evidence, especially from the scaling of nonlocal observables such as EE [12, 13, 14] and disorder operator [11], reveal an anomalous behavior against CFTs even if the partitioning of the lattice is smooth, i.e. no sharp corners on the boundary. The fact that the sign of the observed log-coefficient is consistent with that of the presence of the Goldstone mode, further suggests there exist finite antiferromagnetic moment, i.e. remaining of the Néel order, at the transition point [32]. Such evidence promotes the understanding that the Néel-VBS transition is indeed first order, despite being very weak at the SU(2) limit [39]. In this work, we monitor the behavior of both conventional order parameters and nonlocal observables at such transitions, as the anisotropy $\Delta$ in Eq. (1) is tuned from the Heisenberg limit($\Delta=1$) to easy-plane limit ($\Delta=0$).
@@
Table 1: The critical points obtained from the Binder cumulants of $U_{M}$ and $U_{D}$ as shown in Fig. 2 for different $\Delta=0,0.1,0.3,0.4,0.5$ and 0.6 using similar analysis in Fig. 2. The $q_{c}(\infty)$ of $U_{M}$ and $U_{D}$ are identical within two sigma. The detailed analyses are shown in the Appendix A.
At each $\Delta\in[0,1)$, the phase transition between the antiferromagnetic XY (AFXY) and VBS phase at zero temperature happens when tunning $q=Q/(J+Q)$ from zero to $q_{c}$. To determine the critical points, we perform SSE-QMC simulations [43, 44] on the EPJQ model at $\beta=2L$ for different system sizes $L=16,24,32,48,64,96,128$ with $\Delta=0$, $0.1$, $0.3$, $0.5$ and $0.6$. In the AFXY phase, the order parameter is defined as the sublattice magnetization which breaks the U(1) symmetry, with the $x$ component being written as
@@
where $\{x_{i}$$y_{i}\}$ are coordinates of site $i$ in the lattice. In the simulation with finite system sizes, the expectation value of $\langle M_{x}^{2}\rangle$ can be viewed as the square of the AFXY order parameter, whereas $\langle M_{x}\rangle=0$ at finite sizes.
@@
We thus perform the finite-size scaling (FSS) analysis of $\langle M^{2}_{x}\rangle$ and $\langle D^{2}\rangle$ to determine the transition points at the thermodynamic limit (TDL) [45]. The dimensionless quantity Binder cumulants are widely used in the FSS to extract the critical points. For those two ground states in our study, the Binder cumulants are defined as
@@
with $\nu$ the correlation length exponent and $\omega$ the finite-size correction exponent [28, 45, 46].
@@
Repeating such a procedure, we can also get $q_{c}(\infty)$ for different $\Delta$ in Tab. 1. For a given $\Delta<1$, the $q_{c}(\infty)$ obtained from crossings of different dimensionless quantities agree with each other considering the error as large as two sigma, which can be true for the continuous phase transition as well as the first-order one. However, we should point out that $U_{M}$ of the AFXY order parameter for $\Delta=0.1$ becomes negative close to $q_{c}$ at the largest simulated size $L=128$, as shown in Fig. 2 (a), the negative $U_{M}$ can also be regarded as a signature of the first-order phase transition [47].
We have performed the extrapolation of the ordered moments as in Fig. 2 (d) for all the $\Delta$ values, and the obtained results are summarized in Fig. 1 (b). It is clear that $\langle M^{2}_{x}\rangle$ and $\langle D^{2}\rangle$ are small at $\Delta=0.6$ (they are even smaller at the isotropic limit), but they gradually increases as $\Delta\to 0$. We show all these extrapolations in the Appendix A. The enhancement of the ordered moments as $\Delta\to 0$ along the phase AFXY-VBS phase boundary in Fig. 1 (a) suggests evolution from very weakly first-order to stronger ones. As will be shown below, our non-local EE and disorder operator measurements (with smaller systems sizes compared with order parameters) can also capture such intriguing features. Note that the tiny remaining order parameters may appear to contradict the assumption of our FSS analysis that the system is critical and scale invariant. However, as suggested by our data and a reference by Takahashi et al. [39] for larger system sizes, these weakly first-order transitions exhibit very long correlation lengths, which ensures that the assumption of scale invariance in our FSS analysis holds at the system sizes and precision we consider. It is worth noting that the scaling behavior of binder ratio at evident first order transitions has been discussed in Ref. [48].
@@
where $\langle ij\rangle$ denotes the $J_{1}$ bond and $\langle ij\rangle^{{}^{\prime}}$ denotes the $J_{2}$ bond. This model has a well-established (2+1)d O(3) QCP at $(J_{2}/J_{1})_{c}=1.90951(1)$ [45] separating the Néel phase and a symmetric singlet product phase. As will also shown below, the scaling of disorder operator of this O(3) QCP does not have anomalous log-corrections.